Two point charges reacted with each other to produce 4.32 x 10⁻⁴ J. the magnitude of q is 27 nC with a distance from the point charge of 19 cm. Find the magnitude of the point charge

1
Expert's answer
2022-05-12T13:58:04-0400
"W=k\frac{q_1q_2}{r}"

"q_2=Wr/(kq_1)\\
=4.32*10^{-4}*0.19/(9*10^9*27*10^{-19})\\
=3.38*10^{-7}\:\rm C"